Grade 11General Physics

A potential difference of20 KV is applied across an X-ray tube. Find the minimum wavelength ofX-ray generated.

Rahul Sharma
10 Years agoGrade 11

1 Answer

Ashutosh Mohan Sharma
10 Years ago
f = [1.6 x 10-19 x 2 x 104]/ 6.63 x 10-34 = 4.82x1018Hz

Therefore the wavelengthl(= c/f) is 0.621 x 10-10m = 0.621 nm (compared with some 600 nm for yellow light).
c is the speed of light.